M is the midpoint of QR. Q has coodinates (-3,5) and M has coordinates (7,-9) find the coordinates of R.
Answer by Alan3354(21607)

M is the midpoint of QR. Q has coodinates (-3,5) and M has coordinates (7,-9) find the coordinates of R.
---------------------
The simplest was to do this is to find the differences in X and Y separately, and add them to the midpoint.
x2-x1 = 7 - (-3) = 10
y2-y1 = -9 - 5 = -14
Then add to point M
7+10 = 17
-9-14 = -23
The point R is (17,-23)